OpenStudy (anonymous):

prove that sinx+cosx=cos(x-45)^4

OpenStudy (anonymous):

sorry it is r\[\sqrt{2}\cos(x-45)^4\]

OpenStudy (anonymous):

The power is on the entire cos term correct(not just the angle?) ie. \[\cos^4(x-45)\]

OpenStudy (anonymous):

cos(x-45) = cosxcos45 + sinxsin45 = 1/sqrt2 (cosx+sinx) then sinx + cosx = sqrt2 cos (x-45) don't have any power

OpenStudy (anonymous):

and yeah like @Mathmuse said...

OpenStudy (anonymous):

well i know that but can u prove it with explanation pls

OpenStudy (anonymous):

I explained it lol

OpenStudy (anonymous):

well you has to know first that cos(x-y) = cosxcosy + sinxsiny

OpenStudy (anonymous):

hmm well its cos^4 , u proved it as cos only

OpenStudy (anonymous):

well I think it should be only cos ...I think it's incorrect if it is cos^4

OpenStudy (anonymous):

ok cool :D